# OW-Rep: Open World Object Detection with Instance Representation Learning

### v2

@@ Appendix B Additional Results
B.1 Ablation Studies of Distillation Method
B.2 Ablation Studies of Module under Unknown-Unknown split
B.3 Class Similarity Analysis
B.4 Qualitative Results of Inter-proposal Relationships.
Fig. S2 illustrates the results of inter-proposal relationships on the OWOD split. The embedding of giraffe is used as reference. Our method considers other giraffe as highly similar, sheep as moderately similar, and fire hydrant as dissimilar, which reflects their semantic relationship. This demonstrates that our method can simultaneously detect unknown object and extract semantically rich feature from objects. PROB, on the other hand, considers all the proposals as highly similar since they are not trained to distinguish different unknown objects, but to learn a generalizable concept of object. Both UC-OWOD and RNCDL which learns instance embedding with self-supervision struggles to learn semantic relationship between objects. For more comprehensive comparison of the feature space, refer to Fig. 5.
@@ B.3 Class Similarity Analysis
B.5 Qualitative Results of Open-World Tracking.

### v3

@@ D.3 Class Similarity Analysis
D.4 Quantitative results of Unknown Confusion Metrics
Our method slightly increases confusion metrics, possibly due to the refinement of misclassified unknown boxes. The results are shown in Table S10. Moreover, learning nuanced instance relationships rather than enforcing strict class boundaries may weaken the separation between known and unknown classes.
Appendix E Additional Qualitative Results
Table S10: Results for unknown confusion metrics. WI denotes Wilderness Impact, and A-OSE denotes Absolute Open-Set Error.
[…]
PROB 0.0550 4956 0.0317 5596 0.0165 3122
PROB+URM 0.0642 6821 0.0288 4399 0.0197 4360
PROB+ETM 0.0554 5065 0.0277 3667 0.0169 2808
Ours 0.0621 6056 0.0359 7739 0.0232 5099
E.1 Qualitative Results of Inter-proposal Relationships.
Fig. S4 illustrates the results of inter-proposal relationships on the OWOD split. The embedding of giraffe is used as reference. Our method considers other giraffe as highly similar, sheep as moderately similar, and fire hydrant as dissimilar, which reflects their semantic relationship. This demonstrates that our method can simultaneously detect unknown object and extract semantically rich feature from objects. PROB, on the other hand, considers all the proposals as highly similar since they are not trained to distinguish different unknown objects, but to learn a generalizable concept of object. Both UC-OWOD and RNCDL which learns instance embedding with self-supervision struggles to learn semantic relationship between objects. For more comprehensive comparison of the feature space, refer to Fig. S3.
@@ Appendix E Additional Qualitative Results
E.2 Qualitative Results of Open-World Tracking.
@@
Algorithm 1 Embedding Transfer Module (ETM) within a Training Iteration
Input: Image batch $\{I\}$, Ground truth $\{y\}$, Model $M$, DINOv2 model $M_{DINO}$
Hyperparameters: Margin $\delta$, Kernel bandwidth $\sigma$
$\{B_{pred}\},\{Q_{emb}\}\leftarrow M(\{I\})$ $\triangleright$ Get all predictions and query embeddings
$\{b_{k}\}\leftarrow\text{HungarianMatcher}(\{B_{pred}\},\{y\})$ $\triangleright$ Get known proposals
$\{b_{u}\}\leftarrow\text{Top-k ObjectnesswithRefinement}(\{B_{pred}\})$ $\triangleright$ Get top-k unknown proposals from URM
$B\leftarrow\{b_{k}\}\cup\{b_{u}\}$
$Q_{B}\leftarrow\text{GetCorrespEmbeddings}(Q_{emb},B)$
Step 1: Get Source and Instance Embeddings
$Z\leftarrow\text{MLP}(Q_{B})$
$F\leftarrow M_{DINO}(\{I\})$ $\triangleright$ Extract DINOv2 feature map for the image batch. (A one-time pre-processing step.)
$S\leftarrow\text{AvgPool}(F,B)$ $\triangleright$ Average Pool DINO embeddings from each proposals to obtain source embeddings
Step 2: Compute Pairwise Source Similarity (In practice, implemented using vectorized operations for efficiency.)
$N=|S|$
for $i=1$ to NN, $j=1$ to NN do
  $w_{ij}\leftarrow\exp\left(-\frac{\|s_{i}-s_{j}\|_{2}^{2}}{\sigma}\right)$ $\triangleright$ $\sigma$ : Gaussian kernel, $w_{ij}$ : Pairwise Source Similarity
end for
Step 3: Compute Pairwise Instance Distance
for $i=1$ to NN, $j=1$ to NN do
  $d_{ij}\leftarrow\|z_{i}-z_{j}\|_{2}$ $\triangleright$ $d_{ij}$: Pairwise Euclidean distance
end for
Step 4: Compute Relaxed Contrastive Loss
$\mathcal{L}_{et}=\frac{1}{N}\sum_{i=1}^{N}\sum_{j=1}^{N}\left[\mathbf{w}_{ij}\mathbf{d}_{ij}^{2}+(1-\mathbf{w}_{ij})\left[\delta-\mathbf{d}_{ij}\right]_{+}^{2}\right]$ $\triangleright$ $\delta$ : margin
Return $\mathcal{L}_{et}$